if a circular table is to be made that will have a top covered with material that costs $3.50 per square foot. The covering is to cost no more than $60. What is the maximum radius for the top of the table?
Found 2 solutions by mananth, CharlesG2:
Answer by mananth(16949)

if a circular table is to be made that will have a top covered with material that costs $3.50 per square foot. The covering is to cost no more than $60. What is the maximum radius for the top of the table?
60/3.5 = maximum sq. feet possible. =17.14 sq. feet.
Area of circle = pi*r^2
22/7 * r^2= 17.14
R^2= 17.14*7 /22 = 5.45
R= 2.33 feet.
The radius should not be more than 2.33 feet.
mananth@hotmail.com

Answer by CharlesG2(834)

if a circular table is to be made that will have a top covered with material that costs $3.50 per square foot. The covering is to cost no more than $60. What is the maximum radius for the top of the table?
$60/($3.50/square foot) = $60 * square foot/$3.50
60/3.50 = 6000/350 = 600/35 = 120/7 = 17.142857 square feet approximately maximum for square feet allowed and maximum area of the circular table allowed
pi * r^2 = area of circular table = 17.142857
r^2 = 5.456741 approximately
r = 2.335967 feet approximately maximum radius